triangle ABC, angle B is 5 times angle A and angle C is 14 degrees less than 4 times angle A. Find the size of the angles
Found 2 solutions by oberobic, palanisamy:
Answer by oberobic(2304)

We know the sum of the interior angles for all triangles is 180 = A + B + C.
We also are told:
B = 5A
C = 4A-14
.
A + 5A + 4A-14 = 180
10A = 194
A = 19.4
.
B = 5A = 5(19.4) = 97
C= 4(19.4) - 14 = 63.6
.
We can check by computing the sum to see if it is 180.
19.4 + 97 + 63.6 = ?
19.4 + 97 + 63.6 = 180.
.
So the three angles are:
A = 19.4
B = 97
C = 63.6
.
Done

Answer by palanisamy(496)

Given, in triangle ABC, angle B is 5 times angle A and angle C is 14 degrees less than 4 times angle A
So, B = 5A and C = 4A-14
But we know A+B+C = 180
A+5A+4A-14 = 180
10A = 194
A = 194/10 = 19.4
B = 5A = 5*19.4 = 97
C = 4A-14= 4*19.4-14 = 63.6